# Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors

### v2

@@ V The BRRP Method
Our method takes a single RGBD image and produces reconstructions for each object in the scene. We treat the problem as a Bayesian inference problem over an observation described by negative samples. We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV. We use CLIP [39] to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh. We then use SVGD to optimize for a set of samples over map weights. We can generate predicted reconstructions by taking the expected occupancy over our weights for a given location. Figure 3 shows a visual overview of our method.
@@ V-A Negative Samples as Reconstruction Priors
We want to leverage existing mesh assets as our priors during Bayesian reconstruction. We define our prior as a mixture over different objects, $c_{1},...,c_{C}$. Because there is not a direct way to convert a mesh into a Hilbert map, we instead sample points $\tilde{\mathbf{x}}_{c,1},...,\tilde{\mathbf{x}}_{c,Q}\in\mathbb{R}^{3}$ around each object $c$’s mesh. We refer to these samples as the prior samples. We give them labels $\tilde{y}_{c,1},...,\tilde{y}_{c,1}\in\{-1,1\}$ determined by whether they are outside or inside the mesh. Then we simply define our prior using this data combined with a Gaussian prior over weight norm:
$P(\mathbf{w}|c):=$ $P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})$ (5)
$\propto$ $\exp(\lambda\|w\|^{2})\prod_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i})\right),$ (6)
where BCE is the same as in Equation 1.
In order to enforce pose-invariance, we first register a small stored point cloud of the object to the observed points and then transform the prior samples to this reference frame. In practice we use RANSAC [40] and the FPFH features from [41] to perform registration. In order to also have scale invariance, we do a linear scan over 10 different scales and select the the scale that resulted in the most inlier pairs from the registration.
@@ V-B Retrieval-Augmented Priors for Hilbert Maps
Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV. In order to determine which objects to use, we need to compute $P(c|R)$ from Equation 4. In our case, we use CLIP [39] as a zero-shot classifier for our different objects. For each object in our prior, we store a small textual description of the object. These descriptions are then used as classes for CLIP to classify each segmented object. In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object and feed the cropped images as input into CLIP.
Once we have the probability of each object, we retrieve and register the stored point clouds of the top-k objects. After registration, we retrieve the prior samples corresponding to these objects to define our prior according to Equation 5.
@@ V-C Negative Sampling
We adopt the negative sampling method introduced in [4]. The negative sampling method makes the assumption that all objects are lying on or above a planar surface. We begin by labeling the points segmented to each object as occupied for that object. Next, we perform stratified sampling along each camera ray near each object to recover a set of negatively sampled points, labeled as unoccupied. Then, we use RANSAC over points not segmented to any object to recover the flat surface all objects are resting on. This plane is used to randomly sample points in a sphere underneath each object that are near the object. We also label these points as occupied. Finally, we use grid subsampling from [42] to reduce the number of points and increase uniformity of sampled points. We refer to these points and labels as observed samples and denote them as $\{\mathbf{x}_{i}\}_{i\in[S]},\{y_{i}\}_{i\in[S]}$. The entire negative sampling process can be easily parallelized for efficient computation.
@@ V-D SVGD Reconstruction
Once we have retrieved our prior samples and computed our observed samples, we can perform optimization-based reconstruction with SVGD. Given both sets of samples and our prior definition from Equation 5, we have the following posterior distribution:
$P(\{y_{c,i}\}|\{{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})\sum_{c\in\text{topk}}P(c|R)P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w}),$
taking the log and applying Equation 6 gives us the following objective:
$\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$
$+\sum_{c\in\text{topk}}P(c|R)\ln\left[\sum_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i}))\right)\right]$
$+\lambda\|\mathbf{w}\|^{2}+\text{const.}$.
In practice, we introduce multipliers to each term as hyperparameters during optimization, drop the constant, and use means instead of sums, creating the following objective:
$\frac{\lambda_{3}}{S}\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$ (7)
@@
where $K$ is the number of objects retrieved for the prior. This objective is used as the log of the target distribution, $\ln P(\mathbf{w})$, in Equation 2, where we also adopt the original median kernel suggested in [36]. We also opt to use SVGD in a stochastic manner, where both the observed samples and query samples are mini-batched.
From a non-probabilistic standpoint, one can interpret Equation 7 as the likelihood of the observed data, Equation 8 as the object shape prior, and Equation 9 as a regularization term.
@@ VI Experiments
In this section, we aim to experimentally validate the following claims: (1) BRRP is more robust than deep learning methods; (2) BRRP is more accurate than methods that use uninformative priors; (3) BRRP can capture principled uncertainty. We begin by providing details on the experiments such as the baselines and metrics in Section VI-A, then we introduce results and analyses in Section VI-B
@@ VI-A Experimental Details
BRRP Implementation: We use a set of 50 objects from the YCB dataset [43] to act as the prior for our experiments with BRRP. We implement the method in PyTorch and run the method on an NVIDIA RTX GeForce 2070 GPU.
Baselines: We compare our work against two main baselines, V-PRISM [4] and a version of PointSDF [23] that predicts occupancy and is trained on ShapeNet [44] scenes. V-PRISM is a probabilistic mapping method that uses an uninformative prior. This means that it is robust to novel objects, but doesn’t accurately reconstruct object backsides. We refer to this baseline as V-PRISM. In contrast, PointSDF is a learning-based method. This means it can leverage prior information from mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift. We refer to this baseline as PointSDF. When computing reconstruction meshes with PointSDF, a level set of $\tau=0.3$ is used.
Method ShapeNet Scenes YCB Scenes Objaverse Scenes V-PRISM [4] 0.3092 0.5003 0.4640 PointSDF [23] 0.3600 0.4601 0.3471 BRRP (ours) 0.3124 0.5277 0.4809
Table I: Intersection over union (IoU) on procedurally generated scenes from three different mesh datasets. BRRP uses a YCB prior and PointSDF is trained on ShapeNet scenes.
@@
Procedurally Generated Scenes: We use the generated scenes from [4] to evaluate our method. These scenes are constructed with objects from ShapeNet [44], YCB [43], and Objaverse [45] datasets. There are 100 multi-object scenes for each mesh dataset. Each scene contains up to 10 objects. Some meshes in the Objaverse and ShapeNet scenes did not have correctly rendered textures and were instead rendered as plain white objects. Figure 4 contains two example images of these procedurally generated scenes. We also conduct an experiment on robustness where we perturb instance segmentation of the ShapeNet scenes by 2 pixels and evaluate reconstructions.
We evaluate performance on the procedurally generated scenes with two metrics: intersection over union (IoU) and chamfer distance. We refer readers to [4] for further explanation of these metrics.
Real World Scenes: In order to showcase robustness to real-world noise, we evaluate on real world scenes collected with a Kinect depth camera. In order to obtain instance segmentations, we use Grounded SAM [46] along with some depth filters. We evaluate on these real world scenes qualitatively with images of scene reconstruction and visualizing surface uncertainty.
VI-B Results
@@
As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes. It has the highest IoU improvement from V-PRISM on the YCB scenes. A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap of 0.19 (cm) occuring on the YCB scenes. This makes sense because BRRP uses a subset of the YCB objects as its prior.
We can see this improvement qualitatively in Figure 6. While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a large portion of space that the object doesn’t occupy. BRRP is able to more accurately reconstruct these objects. The clearest example of this is the dark green object in the right-most scene in Figure 6.
Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM. It is the most accurate when evaluated on objects in its prior distribution.
Figure 7: (a) IoU of BRRP and PointSDF on ShapeNet scenes with and without shifted segmentations. Our method is more robust to segmentation shifts. (b) An example of a scene and the corresponding point cloud with shifted segmentation.
@@
While PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 5). BRRP also performed better on mesh datasets that PointSDF was not trained on. In Table I, we can see that on Objaverse scenes, where the objects were novel to both methods, BRRP performed better than PointSDF. When measuring chamfer distance, BRRP outperformed PointSDF on all datasets as shown in Figure 5. These results suggests BRRP is more robust to different object distributions than PointSDF.
Next, we evaluate robustness to slightly incorrect instance segmentations. We take the procedurally generated ShapeNet scenes and shift the segmentation over by 2 pixels. In Figure 7, we compare the IoU of BRRP and PointSDF on the scenes with and without the shift. Our method performs better on the shifted scenes compared to PointSDF.
On the real world scenes in Figure 6, BRRP is qualitatively more robust than PointSDF. PointSDF struggles with the noise associated with real world scenes as well as the novel objects. It tends to predict a spherical object on many non-spherical objects in the real world scenes. Our method on the other hand, is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior.
These results suggest BRRP is more robust than PointSDF. Even though by one metric PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed BRRP performs better.

### v3

@@ V The BRRP Method
Our method takes a single RGBD image and produces reconstructions for each object in the scene. We treat the problem as a Bayesian inference problem over an observation described by negative samples. We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV. We use the pretrained CLIP [radford2021learning] foundation model to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh. We then use SVGD to optimize for a set of samples over map weights. We can generate predicted reconstructions by taking the expected occupancy over our weights for a given location. Figure 4 shows a visual overview of our method.
@@ V-A Negative Samples as Reconstruction Priors
We want to leverage existing mesh assets as our priors during Bayesian reconstruction. We define our prior as a mixture model with different objects serving as each component, $c_{1},...,c_{N_{c}}$. Because there is not a direct way to convert a mesh into a Hilbert map prior, we instead sample points $\tilde{\mathbf{x}}_{c,1},...,\tilde{\mathbf{x}}_{c,Q}\in\mathbb{R}^{3}$ around each object $c$’s mesh. We refer to these samples as the prior samples. We give them labels $\tilde{y}_{c,1},...,\tilde{y}_{c,Q}\in\{-1,1\}$ determined by whether they are outside or inside the mesh. Then we simply define our prior using this data and a Gaussian prior over weight norm:
$\hat{P}(\mathbf{w}|c):=P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})$ (5)
$\hskip 22.76219pt\propto\exp(\lambda\|w\|^{2})\prod_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i})\right).$ (6)
In order to enforce pose-invariance for each prior component, we first register a small stored point cloud of the respective object to the observed points and then transform the prior samples to this reference frame. In practice we use RANSAC [fischler1981random] and the FPFH features from [rusu2009fast] to perform registration. In order to also have scale invariance, we do a linear scan over 10 different scales and select the the scale that resulted in the most inlier pairs from the registration.
@@ V-B Retrieval-Augmented Priors for Hilbert Maps
Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV. In order to determine which objects to use, we need to compute $P(c|R)$ from Equation 4. In our case, we use CLIP [radford2021learning] as a zero-shot classifier for our different objects. For each object in our prior, we store a small textual description of the object. These descriptions are then used as classes for CLIP to classify each segmented object. In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object. One could consider these cropped images to be the $R$ in our $P(c|R)$ formulation. We feed the cropped images as input into CLIP. The resulting distribution over objects predicted by CLIP becomes our $P(c|R)$ for each component.
Once we have the probability of each object, we retrieve and register the stored point clouds of the top-$k$ objects. After registration, we retrieve the prior samples corresponding to these objects to define our top-$k$ component prior approximation according to Equation 5.
@@ V-C Negative Sampling
We adopt the negative sampling method introduced in [wright2024vprism]. The negative sampling method makes the assumption that all objects are lying on or above a planar surface. We begin by labeling the points segmented to each object as occupied for that object. Next, we perform stratified sampling along each camera ray near each object to recover a set of negatively sampled points, labeled as unoccupied. Then, we use RANSAC over points not segmented to any object to recover the flat surface all objects are resting on. This plane is used to randomly sample points in a sphere underneath each object that are near the object. We also label these points as occupied. Finally, we use grid subsampling from [thomas2019learning] to reduce the number of points and increase uniformity of sampled points. We refer to these points and labels as observed samples and denote them as $\{\mathbf{x}_{i}\}_{i\in[S]},\{y_{i}\}_{i\in[S]}$. The entire negative sampling process can be easily parallelized for efficient computation.
@@ V-D SVGD Reconstruction
Once we have retrieved our prior samples and computed our observed samples, we can perform optimization-based reconstruction with SVGD. Given both sets of samples and our prior definition from Equation 5, we have the following posterior distribution,
$P(\{y_{c,i}\}|\{{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})\sum_{c\in\text{topk}}P(c|R)P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w}).$
Taking the log and applying Equation 6 gives the objective:
$\ell=\frac{\lambda_{\text{{1}}}}{S}\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$ (7)
@@
where $K$ is the number of objects retrieved for the prior. For practical reasons, we introduce multipliers to each term as hyperparameters during optimization and use means instead of sums for the first two terms. This objective is used as the log of the target distribution, $\ln P(\mathbf{w})$, in Equation 2, where we also adopt the original median kernel suggested in [liu2016stein]. We also opt to use SVGD in a stochastic manner, where both the observed samples and query samples are mini-batched.
From a non-probabilistic standpoint, one can interpret Equation 7 as the likelihood of the observed data, Equation 8 as the object shape prior, and Equation 9 as regularization.
@@ VI Experiments
In this section, we aim to experimentally validate the following claims: (1) BRRP is more robust than deep learning methods; (2) BRRP is more accurate than methods that use uninformative priors; (3) BRRP can capture principled uncertainty. (4) BRRP’s robustness improves the downstream success of grasping. First, we provide details on the experiments, such as the baselines and metrics in Section VI-A, then we show then show the results and analyses in Section VI-B.
Refer to caption Refer to caption
Figure 6: Results pertaining to the speed of BRRP on an NVIDIA RTX 4070 GPU. Left: Ablation on $k$ for YCB scenes highlights the accuracy-speed tradeoff in retrieval-augmented priors. We use $k=3$ for all other experiments. Right: Timing breakdown of BRRP on a real-world scene with 4 objects.
@@ VI-A Experimental Details
BRRP Implementation: We use a set of 50 objects from the YCB dataset [calli2015ycb] to act as the prior for our experiments with BRRP. We implement the method in PyTorch. We also display results for the speed of running BRRP on an NVIDIA RTX 4070 GPU in Figure 6. We use $k=3$ for all experiments, but note that our method’s performance is fairly robust to the $k$ parameter (Figure 6).
Baselines: We compare our work against two main baselines, V-PRISM [wright2024vprism] and a version of PointSDF [van2020learning] that predicts occupancy and is trained on ShapeNet [chang2015shapenet] scenes. V-PRISM is a probabilistic mapping method that uses an uninformative prior. This means that it is robust to novel objects, but doesn’t accurately reconstruct object backsides. We refer to this baseline as V-PRISM. In contrast, PointSDF is a learning-based method. This means it can leverage prior information from mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift. We train two versions of PointSDF. One is trained on scenes containing a wide variety of Shapenet objects, and another trained only on scenes containing the 50 YCB objects in our BRRP prior. We refer to this baseline as PointSDF, and generally use the ShapeNet version when the training dataset is not mentioned. When reconstructing meshes with PointSDF, a level set of $\tau=0.3$ is used.
Method ShapeNet Scenes (IoU $\uparrow$) YCB Scenes (IoU $\uparrow$) V-PRISM [wright2024vprism] 0.3092 0.5003 PointSDF (ShapeNet) [van2020learning] 0.3600 0.4601 PointSDF (YCB)[van2020learning] 0.2426 0.5088 BRRP (ours) 0.3124 0.5277
Table I: Intersection over union (IoU) on procedurally generated scenes from two different mesh datasets. BRRP uses a YCB prior and there are two different PointSDF versions, trained on ShapeNet and YCB scenes respectively. BRRP outperforms V-PRISM, which has an uniformative prior, whereas BRRP is more robust to distributional shift than PointSDF.
@@
Procedurally Generated Scenes: We use the generated scenes from [wright2024vprism] to evaluate our method. These scenes are constructed with objects from ShapeNet [chang2015shapenet], YCB [calli2015ycb], and Objaverse [deitke2023objaverse] datasets. There are 100 multi-object scenes for each mesh dataset. Each scene contains up to 10 objects. Some meshes in the Objaverse and ShapeNet scenes did not have correctly rendered textures and were instead rendered as plain white objects. Figure 5 contains two example images of these procedurally generated scenes. We use the ground-truth masks for experiments on procedural scenes. We also conduct an experiment on robustness where we perturb the masks of the ShapeNet scenes by two pixels and evaluate reconstructions.
We evaluate performance on the procedurally generated scenes with two metrics: intersection over union (IoU) and chamfer distance. We refer readers to [wright2024vprism] for further explanation of these metrics.
Real World Scenes: In order to showcase robustness to real-world noise, we evaluate on real world scenes collected with a Kinect depth camera. In order to obtain instance segmentations, we use Grounded SAM [ren2024grounded] along with some depth filters. We include timing of performing this segmentation in the real-world timing breakdown in Figure 6. We evaluate on these real world scenes qualitatively with images of scene reconstruction and visualizing surface uncertainty.
Grasping Experiment: To understand the downstream impact of our method, we assess its effectiveness on the problem of dexterous grasping from a fixed RGB-D camera. We place an object on the table, segment it, reconstruct its mesh, plan a grasp trajectory, and execute it open-loop. Figure 9 shows the 10 objects used.
Figure 9: Objects and the robot used to assess the impact of our method on dexterous grasping.
For segmentation, we use Grounded SAM [ren2024grounded] with the prompt object that can be picked up with one hand. Meshes are reconstructed with marching cubes on either BRRP or PointSDF reconstructions. We implement a grasp planner [martin-arxiv-fpte] which takes a segmented point cloud as input and moves the robot to a pregrasp configuration that is collision-free according to the reconstruction while also avoiding obstacles according to the scene reconstruction. The robot then closes the hand and lifts the arm. An attempt is successful if the object is lifted. We use KUKA LBR4 with Allegro hand in the experiment. We place each of the 10 objects across 5 different poses on the table, resulting in 50 executed grasp trajectories for each method. We select stable poses where objects are reachable by the robot, visible to the camera, and remain stationary, while ensuring diversity in positions and orientations. We report the grasp success rate per method across all 50 trajectories.
VI-B Results & Discussion
@@
As showcased in Table I, BRRP outperforms V-PRISM on both sets of procedurally generated scenes. It has the highest IoU improvement from V-PRISM on the YCB scenes. A similar pattern can be seen in the chamfer distances in Figure 7, where BRRP consistently outperforms V-PRISM, with the biggest gap of 0.19 (cm) occuring on the YCB scenes. This makes sense because BRRP uses a subset of the YCB objects as its prior.
We can see this improvement qualitatively in Figure 8. While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a large portion of space that the object doesn’t occupy. BRRP is able to more accurately reconstruct these objects. The clearest example of this is the dark green object in the right-most scene in Figure 8. Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM. It is the most accurate when evaluated on objects in its prior distribution.
Figure 10: (left) IoU of BRRP and PointSDF on ShapeNet scenes with and without shifted segmentations. Our method is more robust to segmentation shifts. (right) An example of a scene and the corresponding point cloud with shifted segmentation.
@@
While the ShapeNet version of PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 7). BRRP also outperformed the YCB version of PointSDF on both YCB and ShapeNet (Table I). Generally, BRRP performed significantly better than each PointSDF version on mesh datasets the specific version was not trained on. When measuring chamfer distance, BRRP outperformed the ShapeNet version of PointSDF on all datasets as shown in Figure 7, including the Objaverse dataset, where objects were out-of-distribution for both methods. These results suggests BRRP is more robust to different object distributions than PointSDF. Next, we evaluate robustness to slightly incorrect image segmentations. We take the procedurally generated ShapeNet scenes and shift the segmentation over by 2 pixels. In Figure 10, we compare the IoU of BRRP and ShapeNet version of PointSDF on the scenes with and without the shift. Our method performs better on the shifted scenes compared to PointSDF.
On the real-world scenes in Figure 8, BRRP is qualitatively more robust than PointSDF. PointSDF struggles with the noise inherent in real world scenes as well as novel objects. It tends to predict a spherical object on many non-spherical objects in the real world scenes. Our method is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior. These results suggest BRRP is more robust than PointSDF. Even though by one metric the ShapeNet-trained PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed, BRRP performs better.
@@
Figure 12: BRRP builds precise representations, which are critical for downstream dexterous manipulation. Here, the water (out-of-distribution) is particularly large, and the dexterous hand needs to arrive at a pre-grasp pose that is very close to the surface of the bottle, all the while requiring the fingers to avoid colliding.
@@
Insight 4: BRRP’s robust reconstructions improve the downstream success of dexterous grasping.
The success rate of BRRP and the PointSDF baseline is 64% and 40% respectively on the downstream grasping experiment. The difference in success rates stems from the modeling of the obstacles, consisting of the reconstructed object mesh and the table (fixed). The collision modeling directly impacts how close the robot can get to the object before executing the grasp. This directly affects whether the object will get knocked over or bumped into as the gripper moves to the pre-grasp configuration. BRRP exhibits higher accuracy and robustness in reconstruction, which creates an over 50% increase in grasp success rates. Figure 12 illustrates an example and highlights the impact of BRRP in significantly improving downstream grasp success. While the difference in BRRP and PointSDF grasp success can be easily attributed to reconstruction quality, we hypothesize the 36% of grasp failures with BRRP stem from a mixture of reconstruction, grasp selection, and motion planning inaccuracies.
Limitations: As shown in Figure 6, the registration portion of our method grows with the number of objects, which can be resolved by fast learning-based registration methods. While our method is robust to viewpoint, occlusion, and scale; it is still limited by the quality of segmentation, depth accuracy, and its prior database. Advances in segmentation methods will further improve BRRP performance. While BRRP can robustly handle OOD data, the reconstruction and uncertainty remain marginally biased towards the in-distribution dataset, which is common for data-driven approaches. This can be seen in Figure 2 and Figure 8, and can be mitigated with a larger, more diverse prior database.